# Weak measurement-based protocol for ergotropy protection in open quantum batteries

### v1

@@ III.3 Total ergotropy gain and success probability
The total ergotropy changes induced during the protocol, composed of both the incoherent and coherent components, are depicted in Fig. 9 for the same parameters introduced previously. Note that, in this case, the resulting battery discharging process after the weak measurement (dashed blue line) is slower than the one obtained without performing it (continuous red line). Such a difference in ergotropy loss leads to the crossing of both curves and, ultimately, more ergotropy is saved after some dissipation time. Nevertheless, despite satisfying the constrain $\varepsilon_{m\tilde{w}}(\tau_{\gamma})=\mathcal{W}_{m\tilde{w}}(\tau_{\gamma})=0$, the reversal measurement gives even more ergotropy to the battery, providing a positive gain at the end of the protocol of ¯$\overline{\mathcal{R}}_{mw}(\tau_{\gamma})\approx 0.76\times 10^{-1}$. The probability of success for this particular example is $\Pi_{m\tilde{w}}(\tau_{\gamma})\approx 0.57$.
IV Discussion
@@
Given recent advancements in quantum thermodynamics [100] and the growing ability to manipulate quantum systems [77], the development of quantum technologies for energy control and management has become increasingly crucial [101, 102, 103, 104]. In this context, the design and analysis of efficient quantum devices for energy storage have gained significant interest, with proposals across different physical platforms [25, 26, 27, 23, 28, 29, 30, 5, 31, 32]. Nevertheless, due to the fragile nature of quantum systems, realistic implementations must consider environmental interactions, which often lead to unwanted discharging. Developing methods to counteract such discharge effect has thus become a key focus for practical quantum battery applications, and this work contributes to this ongoing effort.
@@
Here, we introduced a protocol for mitigating the inevitable discharging process in a single qubit open quantum battery by applying two selective weak measurements. More specifically, a first weak measurement is performed right after the battery charging -before being subjected to environmental interaction-, followed by a reversal measurement after some dissipation time. The protocol’s operating points were set by thermodynamic constrains, namely by simultaneously meeting the conditions $\varepsilon_{mw}(\tau)=0$ and $\mathcal{W}_{mw}(\tau)=0$, which respectively define the proper parameter space for minimum external energetic and ergotropic influences. Such conditions can be achieved due to the symmetry of the weak measurement and its reversal. In this way, we showed that by properly executing the protocol, one can obtain positive gain, i.e., the final ergotropy can be higher compared to a purely discharged quantum battery, effectively mitigating the effects of environmental interactions. Nevertheless, it is worth mentioning the previous constraints can be relaxed, provided the ergotropy shift $\mathcal{W}_{mw}(\tau)$ remains below the protocol-induced ergotropy gain; in other words, even allowing for small extra recharging from the measurements can be advantageous, depending on the success probability. In the more restrictive case—where both measurements result in zero energy and ergotropy shifts—we analyzed the differences in behavior for the incoherent and coherent components of ergotropy.
@@
To demonstrate our method, we selected specific parameter examples. Since any quantum system decays over the time scale determined by the dissipation rate into the environment, we concentrated on the energy (ergotropy) mitigation in the period of $\tau_{\gamma}=1/\gamma$. We observed that achieving higher ergotropy gains comes at the cost of lower probabilities of success, and vice versa, indicating an inverse relationship. Notably, the success probability landscape of the protocol remains invariant with respect to the initial coherence. Through our examples, we demonstrated how the coherent ergotropy gain can positively contribute at higher probabilities, showcasing how quantum features like coherence can be effectively harnessed to enhance the protocol’s overall performance.
@@
The results presented here can be straightforwardly expanded to consider a protocol for multi-qubit batteries, exploring whether the obtained gain scale with an increasing number of systems (cells). This would involve considering the role played by the internal coupling between different cells of the battery and determining how interactions among qubits might affect overall performance. Additionally, it would be valuable to understand how quantum correlations such as entanglement and quantum discord [105, 106] behave within this weak measurement-based protocol, especially as it could impact both the stability and efficiency of the quantum battery. Examining the energetic cost of measurements, following Landauer’s principle, would also be important, as this would provide insights into the thermodynamic cost associated with implementing the weak measurements, allowing for a more comprehensive assessment of the protocol’s practicality.
@@
Finally, optimal parameter choices for successfully applying the protocol will likely depend on system-specific and environmental constraints. In realistic setups, it may not always be possible to control factors like dissipation time, temperature or to precisely modulate measurement strength, which could limit the achievable gains. Future work should consider these constraints and investigate adaptive protocols that can respond to variable conditions, enabling the practical application of this approach across a wider range of quantum battery designs and operational environments.
V Acknowledgements

### v2

@@ III.3 Total ergotropy gain and success probability
The total ergotropy changes induced during the protocol, composed of both the incoherent and coherent components, are depicted in Fig. 10 for the same parameters introduced previously. Note that, in this case, the resulting battery discharging process after the weak measurement (dashed blue line) is indeed slower than the one obtained without performing it (continuous red line). Such a difference in ergotropy loss leads to the crossing of both curves and, ultimately, more ergotropy is saved after some dissipation time. Moreover, despite satisfying the constrain $\varepsilon_{m\tilde{w}}(\tau_{\gamma})=\mathcal{W}_{m\tilde{w}}(\tau_{\gamma})=0$, the reversal measurement gives even more ergotropy to the battery, providing a positive gain at the end of the protocol of ¯$\overline{\mathcal{R}}_{mw}(\tau_{\gamma})\approx 0.76\omega\times 10^{-1}$, or $100\times\overline{\mathcal{R}}_{mw}(\tau_{\gamma})/\mathcal{R}_{i}\approx 8.57\%$ of saved charge with the protocol. The probability of success for this particular example is $\Pi_{m\tilde{w}}(\tau_{\gamma})\approx 0.57$.
IV Extension to multi-qubit open quantum batteries
Practical quantum batteries will likely require systems composed of multiple energy-storing units, or battery cells. Here, we extend our two-time weak measurement protocol to quantum batteries with $N$-qubit cells. To demonstrate this approach, we analyze a two-qubit interacting system and examine its behavior when quantum correlations are present. For a quantum battery comprising $N$ identical qubits, the Hamiltonian takes the form $\hat{H}=\omega\sum_{k}^{N}|e\rangle\langle e|_{k}+\sum_{k\neq j}^{N}J_{kj}\hat{\sigma}_{+}^{k}\hat{\sigma}_{-}^{j}$, where $J_{kj}$ represents the coupling strength between the $k$-th and $j$-th cells, determined by the specific architecture. The open battery dynamics are strongly influenced by parameter ranges and underlying assumptions. We adopt a local master equation derived under standard Born-Markov approximations with a common environment [59, 146]. This framework requires that internal couplings satisfy $J_{kj}\ll\omega$ for all $k,j$, allowing jump operators to be computed in their local bare Hamiltonian basis. Under these conditions, the time evolution follows the structure of Eq. (3), with dissipators given by:
$\begin{split}\mathcal{D}_{\downarrow}[\bm{\cdot}]&=\gamma(1-f)\left(\hat{S}_{+}\bm{\cdot}\hat{S}_{-}-\frac{1}{2}\left\{\hat{S}_{-}\hat{S}_{+},\bm{\cdot}\right\}\right),\\ \mathcal{D}_{\uparrow}[\bm{\cdot}]&=\gamma f\left(\hat{S}_{-}\bm{\cdot}\hat{S}_{+}-\frac{1}{2}\left\{\hat{S}_{+}\hat{S}_{-},\bm{\cdot}\right\}\right),\end{split}$ (37)
where $\hat{S}_{\pm}=\sum_{k}^{N}\hat{\sigma}_{\pm}^{k}$ represent the raising and lowering ladder operators. This dynamics can be realized in atomic chains [147, 148], for instance. Notably, as a consequence of the local approach, the expression above is independent of internal couplings $J_{kj}$. Nevertheless, their effects are incorporated through the Hamiltonian component of the dynamics. It is important to note that Eq. (37) includes not only local dissipators but also cross-terms that couple different battery cells. Considering two qubits, for instance, these contributions may arise when the full secular approximation fails for resonant qubits, resulting in environment-induced coupling [146]. Common environments have been shown to produce significant effects, including collective phenomena [149, 76, 150] and steady-state entanglement [151, 152, 153], with various potential applications [154, 155, 156, 157, 158]. Also, it is important to highlight the dynamics induced by this master equation and, therefore, the stationary state strongly depends on the choice of the initial state.
For a $N$-cell battery, the TWM protocol follows a direct generalization of the procedure introduced in Sec. II. The protocol involves applying multiple simultaneous local weak measurements during both the measurement and reversal steps. Specifically, for each $k$-th cell, measurements $\hat{M}_{k}$ and $\hat{W}_{k}$ are applied with respective forces $m_{k}$ and $w_{k}$, such that: $\hat{M}=\bigotimes_{k}^{N}\hat{M}_{k}$ and $\hat{W}=\bigotimes_{k}^{N}\hat{W}_{k}$, where
$\begin{split}\hat{M}_{k}&=|g\rangle\langle g|_{k}+\sqrt{1-m_{k}}|e\rangle\langle e|_{k},\\ \hat{W}_{k}&=\sqrt{1-w_{k}}|g\rangle\langle g|_{k}+|e\rangle\langle e|_{k}.\end{split}$ (38)
The operational points of the protocol are defined by regions in the parameter space that ensure null energetic and ergotropic shifts, i.e., $\varepsilon(\tau)=0$ and $\mathcal{W}(\tau)=0$, respectively. While the number of measurement strengths increases with the number of cells, s.t. $\textbf{m}=(m_{1},...,m_{N})$ and $\textbf{w}=(w_{1},...,w_{N})$, operational points remain achievable by relying on numerical methods. Furthermore, it is important to note that not all local measurements are required to obtain some gain, as long as the net effects of all local weak and reversal measurements compensate for each other. This implies that not all cells are required to be measured for a successful protocol execution.
To demonstrate this procedure, we will examine a two-qubit quantum battery, focusing on the feasibility of our protocol when quantum correlations, such as entanglement and quantum discord, are present. Without loss of generality, the following analysis will assume $\gamma=10^{-2}$ and $J=2\omega\gamma$, s.t. $J\ll\omega$.
Figure 11: $X$-state properties. (a) Ergotropy and (b) concurrence in terms of $1-q$, assuming $J<\omega$. The symmetry of the ergotropy reveals parameter choices where systems share the same ergotropy but differ in the presence or absence of correlations.
Figure 12: Ergotropy and concurrence dynamics for the $X$-state. (a) Ergotropy and (b) concurrence for $\omega=1$. The solid red and dashed blue curves represent $q=0.1$ and $q=0.9$, respectively. While the ergotropy decays rapidly in both cases, the concurrence exhibits different behaviors: for $q=0.1$, it initially increases due to the dissipation dynamics given by Eq. (37) before eventually decreasing, whereas for $q=0.9$, it decreases monotonically throughout the process. It was assumed $f=0.3$, $\gamma=10^{-2}$ and $J=2\omega\gamma$.
IV.1 Example: Two-qubit open quantum battery
Given two uncoupled qubit batteries individually prepared in the same incoherent state $\hat{\rho}_{0}=q|g\rangle\langle g|+(1-q)|e\rangle\langle e|$, one can unitarily produce an $X$-state in the form of
$\hat{\rho}_{X}(q)=\left(\begin{array}[]{cccc}\rho_{11}&0&0&\rho_{14}\\ 0&\rho_{22}&\rho_{23}&0\\ 0&\rho_{23}^{*}&\rho_{33}&0\\ \rho_{14}^{*}&0&0&\rho_{44}\end{array}\right),$ (39)
that optimally enhances the quantum correlations between the cells. The resulting bipartite state $\hat{\rho}_{X}(q)$ is given by $\rho_{11}=\rho_{44}=q/2$, $\rho_{14}=q^{2}-q/2$, $\rho_{23}=0$, $\rho_{22}=q(1-q)$, and $\rho_{33}=(1-q)^{2}$ [159, 160] (see Appendix C for more details). Bipartite local weak measurements and their reversal have been shown to effectively restore quantum entanglement lost due to decoherence [124] and dissipation [139]. Importantly, the symmetries of this model allow a meaningful comparison between two distinct scenarios with the same initial ergotropy—one where entanglement is present and another where it is absent due to entanglement sudden death.
Figure 13: Energy shift profile at $\tau_{\gamma}$ in terms of the reversal measurement strengths $w_{1}$ and $w_{2}$ for (a) $q=0.1$ with weak measurement strengths $\textbf{m}\approx(0.5,0.6)$ and (b) $q=0.9$ with weak measurement strengths $\textbf{m}=(0.5,0.9)$. The red stars highlight the operational points where $\varepsilon(\tau_{\gamma})=0$ (dashed blue curve) and $\mathcal{W}(\tau_{\gamma})=0$ (continuous red curve) are simultaneously satisfied. It was assumed $f=0.3$, $\gamma=10^{-2}$, $J=2\omega\gamma$ and $\omega=1$.
Since $J<\omega$ by hypothesis, the ergotropy of $\hat{\rho}_{X}(q)$ is given by $\mathcal{R}\left[\hat{\rho}_{X}(q)\right]=\omega\left|1-2q\right|$, while the entanglement, quantified by the concurrence, is $\mathcal{C}[\hat{\rho}_{X}(q)]=\max\left[\left\{0,2q^{2}-q-2(1-q)\sqrt{q(1-q)}\right\}\right]$. Figure 11 illustrates the behavior of these quantities as functions of the excited population $1-q$. In Fig. 11(a), the ergotropy exhibits a symmetric profile with respect to $q$, such that pairs of different $X$-states share the same ergotropy: $\mathcal{R}[\hat{\rho}_{X}(1/2-x)]=\mathcal{R}[\hat{\rho}_{X}(1/2+x)]=2\omega|x|$, for $x\in[-1/2,1/2]$. Meanwhile, Fig. 11(b) reveals entanglement sudden death, with concurrence vanishing for $q\leq 0.7$.
Figure 14: Ergotropy comparison without initial entanglement. Ergotropy changes with (dashed blue) and without (continuous red) the protocol for (a) total ergotropy, (b) incoherent ergotropy, and (c) coherent ergotropy assuming $q=0.1$, $m_{1}=0.5$, $m_{2}\approx 0.6$ and $w_{1}=w_{2}\approx 0.21$. For this particular state, which initially has no entanglement, the protocol is more effective in recovering the incoherent part of the ergotropy.
Figure 15: Ergotropy comparison with initial entanglement. Ergotropy changes with (dashed blue) and without (continuous red) the protocol for (a) total ergotropy, (b) incoherent ergotropy, and (c) coherent ergotropy assuming $q=0.9$, $m_{1}=0.5$, $m_{2}=0.9$, $w_{1}\approx 0.97$ and $w_{2}=0.17$. For the last step, it was assumed $\omega=1$. For the initially entangled state, the protocol performs effectively in recovering both the incoherent and coherent parts of the ergotropy, demonstrating its feasibility when quantum correlations are present.
In addition, assuming the initial charged state $\hat{\rho}_{X}(q)$, the battery dynamics under the master equation given by Eq. (37) can be shown to maintain the $X$ structure of Eq. (39) invariant, such that
$\hat{\rho}(t)=\left(\begin{array}[]{cccc}\rho_{11}(t)&0&0&\rho_{14}(t)\\ 0&\rho_{22}(t)&\rho_{23}(t)&0\\ 0&\rho_{23}^{*}(t)&\rho_{33}(t)&0\\ \rho_{14}^{*}(t)&0&0&\rho_{44}(t)\end{array}\right).$ (40)
Due to ergotropy symmetry, distinct initial states can share the same ergotropy yet evolve differently, leading to separate steady states. Interestingly, certain initial conditions enable the generation of entanglement and/or an increase in ergotropy during both transient dynamics and in the steady state. Figure 12 depicts the time evolution of (a) ergotropy and (b) concurrence for $q=0.1$ and $q=0.9$ over the relevant protocol window $\gamma t\in[0,1]$. Initially, $\mathcal{R}[\hat{\rho}_{X}(0.1)]=\mathcal{R}[\hat{\rho}_{X}(0.9)]=0.8\omega$, while the concurrence values differ: $\mathcal{C}[\hat{\rho}_{X}(0.1)]=0$ and $\mathcal{C}[\hat{\rho}_{X}(0.9)]=0.66$, for $\omega=1$. The ergotropy decays monotonically in both cases, but the concurrence evolves differently. For $q=0.1$, the system generates correlations, leading to an initial increase in concurrence, followed by a decrease and slight oscillations. In contrast, for $q=0.9$, the concurrence monotonically decreases due to environmental interactions.
Figure 16: Ergotropy gain $\times$ success probability in terms of $q$ and $m_{2}$, $w_{1}$ and $w_{2}$, computed at $\tau_{\gamma}$, $\omega=1$ and $m_{1}=0.5$ for (a) $f=0.30$ (b) $f=0.15$. The red spots highlight the operational points, with zero ergotropy shift. The orange dots (vertical lines) at $q=0.1$ and $q=0.9$ in Fig. (a) highlight the operational points used for Figs. 14 and 15, respectively.
Given the number of parameters involved, particularly the sets $\textbf{m}=(m_{1},m_{2})$ and $\textbf{w}=(w_{1},w_{2})$, determining the exact analog of Eq. (24) for the null energetic shift surface is challenging. Nevertheless, as previously discussed, operational points can still be identified. To illustrate this, Fig. 13 presents the energetic shift $\varepsilon(\tau_{\gamma})/\omega$ profiles for $\omega=1$ as a function of the reversal measurement strengths w. Specifically, panel (a) corresponds to $q=0.1$ with weak measurement strengths $\textbf{m}\approx(0.5,0.6)$, while panel (b) shows the case for $q=0.9$ with $\textbf{m}=(0.5,0.9)$. In each plot, the dashed blue and continuous red curves indicate the points where $\varepsilon(\tau_{\gamma})=0$ and $\mathcal{W}(\tau_{\gamma})=0$, respectively. The intersections of these curves, marked by red stars, correspond to the operational points for the selected parameters and used for the following analysis.
Figures 14 and 15 confirm the effectiveness of the TWM protocol for two-qubit batteries in both initially non-entangled and entangled scenarios, demonstrating its potential for scaling to more complex coupled systems. In Fig. 14, the (a) total, (b) incoherent, and (c) coherent ergotropy evolutions during the protocol (dashed blue curves) are compared to purely dissipative dynamics (continuous red curves) for $q=0.1$, with measurement strengths $\mathbf{m}\approx(0.5,0.6)$ and $\mathbf{w}\approx(0.21,0.21)$. Across all ergotropy components, the weak measurement initially reduces the energy, followed by a monotonic decay, and a subsequent increase due to the reversal measurement. At the end of the protocol, a net ergotropy gain of ¯$\overline{\mathcal{R}}(\tau_{\gamma})\approx 0.49\omega\times 10^{-1}$ is achieved, with a success probability of $\Pi(\tau_{\gamma})\approx 0.37$. This represents a $100\times\overline{\mathcal{R}}(\tau_{\gamma})/\mathcal{R}_{i}\approx 6.12\%$ of saved charge due to the protocol execution. Similarly, Fig. 15 examines the same quantities for initially quantum correlated batteries with $q=0.9$, using measurement strengths $\mathbf{m}=(0.5,0.9)$ and $\mathbf{w}\approx(0.97,0.17)$ and $\omega=1$. Despite having the same initial charge as in the $q=0.1$ case, the ergotropy evolution differs significantly. The strong weak measurement removes almost all ergotropy. Furthermore, due to collective dissipation in the shared reservoir, some initial states experience an ergotropy increase. Here, while the coherent ergotropy decreases, the incoherent and total ergotropy exhibit some growth. The overall ergotropy gain is ¯$\overline{\mathcal{R}}(\tau_{\gamma})\approx 6.1\omega\times 10^{-1}$, with a success probability of $\Pi(\tau_{\gamma})\approx 0.09$. Interestingly, due to environment-assisted charging, the final ergotropy after the reversal measurement slightly exceeds the initial value, yielding $\mathcal{R}_{iv}-\mathcal{R}_{i}\approx 0.3\omega\times 10^{-2}$ and a recover of $100\times\overline{\mathcal{R}}(\tau_{\gamma})/\mathcal{R}_{i}\approx 75.81\%$ of the initial charge. The protocol also affects the quantum correlations within the cells (see Appendix C). In particular, for $q=0.1$, the TWM enhances the final concurrence compared to the purely dissipative process, with gain $\overline{\mathcal{C}}=\mathcal{C}[\hat{\rho}_{mw}(\tau_{\gamma})]-\mathcal{C}[\hat{\rho}(\tau_{\gamma})]>0$.
Finally, Fig. 16 illustrates the relation between ergotropy gain ¯$\overline{\mathcal{R}}(\tau_{\gamma})$, success probability $\Pi(\tau_{\gamma})$, and $q$, evaluated at the operational time $\tau_{\gamma}$, for fixed measurement strength $m_{1}=0.5$, and temperature parameter (a) $f=0.30$, (b) $f=0.15$. All displayed configurations satisfy the zero-energy shift condition. The red lines highlight representative operational pathways with zero ergotropy shift $\mathcal{W}(\tau_{\gamma})=0$, while the vertical dashed orange lines in the $q=0.1$ and $q=0.9$ surfaces in (a) mark the configurations analyzed in Figs. 14 and 15, respectively. From both panels of Fig. 16, we observe that temperature variations have a more pronounced effect on higher values of $q$. Additionally, for lower $q$, the operational points are predominantly confined to the low-gain, high-success-probability regime, whereas higher $q$ values exhibit the opposite behavior, favoring high-gain but lower success probabilities. In particular, for $q\lesssim 0.7$—where the initial $X$-states are non-entangled [cf. Fig. 11(b)]—a clear trade-off emerges: achieving higher success probabilities $\Pi(\tau_{\gamma})$ typically comes at the cost of lower ergotropy gains ¯$\overline{\mathcal{R}}(\tau_{\gamma})$. This is visible in the downward-sloping red trajectories embedded within the green and black surfaces in that regime. In contrast, for $q\gtrsim 0.7$, where the initial states exhibit entanglement, this trade-off becomes less stringent—several operational points simultaneously offer both appreciable success probabilities and enhanced ergotropy gain. This indicates that initial entanglement may play a constructive role in alleviating the gain–probability trade-off.
Therefore, the TWM protocol can be effectively applied to multi-qubit quantum batteries. Specifically, we compared its behavior for two-qubit batteries in a common reservoir, where the initial states had the same charge but differed in their quantum correlations. Notably, due to collective dissipation, the initially correlated cells not only mitigated energy loss but even gained additional charge, with the cost of it being unlikely. Finally, we note that similar behavior and conclusions are expected to hold for other values of the internal coupling $J$, provided that $J\ll\omega$. In the local approach, the coupling primarily affects the unitary dynamics. Stronger couplings, however, fall outside the validity of Eq. (37), requiring a global master equation for an accurate description.
V Discussion
@@
Given recent advancements in quantum thermodynamics [161] and the growing ability to manipulate quantum systems [108], the development of quantum technologies for energy control and management has become increasingly crucial [162, 163, 164, 165]. In this context, the design and analysis of efficient quantum devices for energy storage have gained significant interest, with proposals across different settings and physical platforms [47, 48, 49, 46, 53, 54, 56, 4, 57, 58, 52]. Nevertheless, due to the fragile nature of quantum systems, realistic implementations must consider environmental interactions, which often lead to unwanted discharging. Developing methods to counteract such discharge effects has thus become a key focus for practical quantum battery applications [80], and this work contributes to this ongoing effort.
@@
Here, we introduced a protocol for mitigating the inevitable discharging process in open quantum batteries by applying two selective weak measurements. More specifically, a first weak measurement is performed right after the battery charging—before being subjected to environmental interaction—, followed by a reversal measurement after some dissipation time. The protocol’s operating points were set by thermodynamic constraints, namely by simultaneously meeting the conditions $\varepsilon_{mw}(\tau)=0$ and $\mathcal{W}_{mw}(\tau)=0$, which respectively define the proper parameter space for minimum external energetic and ergotropic influences. Such conditions can be achieved due to the symmetry of the weak measurement and its reversal. In this way, we showed that by properly executing the protocol, one can obtain positive gain, i.e., the final ergotropy can be higher compared to a purely discharged quantum battery, effectively mitigating the effects of environmental interactions. Nevertheless, it is worth mentioning the previous constraints can be relaxed, provided the ergotropy shift $\mathcal{W}_{mw}(\tau)$ remains below the protocol-induced ergotropy gain; in other words, even allowing for small extra recharging from the measurements can be advantageous, depending on the success probability. In the more restrictive case—where both measurements result in zero energy and ergotropy shifts—we analyzed the differences in behavior for the incoherent and coherent components of ergotropy.
@@
To validate the TWM protocol, we selected specific parameter examples. Since any quantum system decays over the time scale determined by the dissipation rate into the environment, we concentrated on the energy (ergotropy) mitigation in the period of $\tau_{\gamma}=1/\gamma$. It is also worth mentioning that the role of temperature, encoded in $f$, directly influences the dissipation process, affecting both the stationary state and the rate at which it is reached. Nevertheless, the choice of temperature does not fundamentally constrain the presented results. The findings are expected to hold for different temperature values, provided that the relevant parameters—particularly the weak and reversal measurement strengths, and dissipation time—are appropriately adapted. In general, we observed that achieving higher ergotropy gains comes at the cost of lower probabilities of success, and vice versa, indicating an inverse relationship. Notably, the success probability landscape of the protocol for a single qubit battery remains invariant with respect to the initial coherence. Through our examples, we demonstrated how the coherent ergotropy gain can positively contribute at higher probabilities, showcasing how quantum features like coherence can be effectively harnessed to enhance the protocol’s overall performance.
Moreover, we discussed how to extend the analysis to multi-cell quantum batteries, demonstrating how the two-time weak measurement protocol generalizes to systems composed of $N$ qubits. To validate this approach, we examined a specific scenario consisting of two interacting qubits sharing a common environment. This kind of dynamics captures non-trivial phenomena due to collective dissipation, such as entanglement and ergotropy increase. To contemplate the potential role of correlations, we considered $X$-states as a case study. In particular, we employed the symmetries within these states to compare two distinct scenarios with the same initial ergotropy, one not containing entanglement and the other containing it. The protocol is shown to work in both cases, mitigating the ergotropy loss compared to the purely dissipative case.
@@
The results presented here can be straightforwardly expanded to consider multi-qubit batteries in different relevant regimes. In particular, future work should consider strongly coupled cells and characterize the role of internal coupling on the overall performance of the TWM protocol. We highlight that this regime is beyond the validity of Eq. (37), requiring the use of a global master equation for an accurate dynamic description and thorough analysis of the coupling. Additionally, it would be valuable to deeply understand the general features of quantum correlations such as entanglement and quantum discord [166, 167] within the overall performance of the TWM protocol, especially how it could scale with the number of cells and depend on the configuration of internal couplings. Besides, it would be relevant to consider how the recently proposed notions of parallel ergotropy [168], local ergotropy [169, 170], and extended local ergotropy [171] change during such a procedure. Investigating its viability for strong system-environment couplings [170] and non-Markovian environments [37, 66] is also of fundamental interest. Examining the energetic cost of measurements, following Landauer’s principle [172], would also be important, as this would provide insights into the thermodynamic cost associated with implementing the weak measurements, allowing for a more comprehensive assessment of the protocol’s practicality.
@@
Finally, optimal parameter choices for successfully applying the protocol will likely depend on system-specific and environmental constraints. In realistic setups, it may not always be possible to control factors like dissipation time and temperature or to precisely modulate measurement strength, which could limit the achievable gains. Future work should examine optimization techniques and consider these constraints to investigate adaptive protocols that can respond to variable conditions, enabling the practical application of this approach across a wider range of quantum battery designs and operational environments. In this context, it is worth highlighting that the proposed protocol is practically feasible using state-of-the-art platforms. The implementation of weak measurements and measurement reversal operations for mitigating decoherence has already been demonstrated [124, 125, 137, 121], while quantum batteries have been realized in comparable settings. Thus, by considering the thermodynamic constraints, one could in principle achieve the protocol’s working regime. Considering recent experimental advancements in both areas, NMR [58, 109] and superconducting circuits [46, 52] emerge as promising candidates for the application of the two-time weak measurement (TWM) protocol introduced in this work.
VI Acknowledgements